# Entanglement Through Topological Defects: Reconciling Theory with Numerics

### v1

@@ Abstract
Present theoretical predictions for the entanglement entropy through topological defects are violated by numerical simulations. In order to resolve this, we introduce a paradigm shift in the preparation of reduced density matrices in the presence of topological defects, and emphasize the role of defect networks with which they can be dressed. We consider the cases of grouplike and duality defects in detail for the Ising model, and match all numerically found entanglement entropies. Since our construction functions at the level of reduced density matrices, it accounts for topological defects beyond the entanglement entropy to other entanglement measures.
1. Introduction.— Defects and interfaces appear ubiquitously in nature, from impurities in materials AffleckGFactor over quantum dots sierra2014strongly all the way to D-branes in string theory Polchinski:1998rr. The last decade has, moreover, witnessed a reformulation of our understanding of symmetries to what is now known as generalized symmetry Shao:2023gho – a development driven mainly by topological defects. These are completely energy-transmissive and can therefore be deformed at will, until fields or boundaries are encountered, where topological defects act in various interesting ways. In addition to implementing conventional group transformations, topological defects can implement, for instance, order-disorder dualities Frohlich_2007, most famously Kramers-Wannier duality in the Ising model Frohlich:2004ef.
@@
The effect of topological defects, in contrast, on entanglement is more nuanced. They appear for instance naturally as tool for symmetry resolution of entanglement spectra Belin_2013; Goldstein:2017bua; Zhao_2021; Weisenberger_2021; Zhao:2022wnp; northe2023entanglement; Kusuki:2023bsp; Choi:2024wfm; Kusuki:2024gss; Heymann_2025; Northe:2025qcv; benini2025entanglementasymmetryhighernoninvertible; Bhattacharyya:2025tmg, where the defects are imposed manually. Situations where the defect forms part of the physical system were studied early on Sakai_2008; Brehm:2015lja; Brehm_2016; Brehm:2017kuz; gutperle2024noteentanglemententropytopological; Gutperle_2016 in high-energy theory. These works predict a subleading contribution to the EE, which, however, have recently been seen to be violated in numerical simulations of the Ising model Roy_2022; Rogerson:2022yim, the 3-state Potts model Sinha:2023hum and the Luttinger liquid Roy:2023wer.
Subleading contributions encode the intricate interplay between boundaries, defects, but also field excitations Ibáñez_Berganza_2012. It is therefore important to have a formalism capable of predicting the effect of topological defects on quantum correlations correctly. It is the purpose of this letter to fill this gap.
It is useful to pinpoint where the theoretical framework of Sakai_2008; Brehm:2015lja; Brehm_2016; Brehm:2017kuz; gutperle2024noteentanglemententropytopological; Gutperle_2016 has deficiencies, so we sketch it. Consider an interface $L$ separating the right-half from the left-half plane, and denominate the positive real line as entangling interval $A$. An reduced density matrix (RDM) $\rho_{A}^{old}$ is prepared Sakai_2008; Brehm:2015lja; Brehm_2016; Gutperle_2016 by identifying two concentric circles centered at the entangling edge $\partial A$, one with radius $\epsilon\ll 1$, and one with radius $1/\epsilon$,
@@
In this way the interval $A$ is compactified onto a circle, and the RDM is a defect-dressed cylinder in which entanglement time runs from bottom to top. Because the identification splits the defect in two, it acts twice on the subregion Hilbert space $\mathcal{H}_{A}$. Stacking $n$ such RDMs and tracing over $\mathcal{H}_{A}$ leads to a torus. Two problems are apparent:
@@
Below, we construct RDMs for an interval $A$, which do not suffer from ((P1)) & ((P2)). Besides having $L$ pierce the entangling edge $\partial A$, we also discuss the situations where $L$ crosses inside and outside of $A$. Our guiding principle is to consider proper choices of defect network that complete the defect piercing the constant time slice. Ultimately, this leads us to introduce twist fields and study their quantum correlations. We work directly at the level of RDMs, so that our formalism’s utility extends beyond the EE to the investigation of other entanglement measures.
After briefly recapitulating essential properties of boundaries and defects, we construct our RDMs and subsequently match their Rényi entropies to numerically obtained EEs in the Ising model Roy_2022.
@@
Note that in the second picture the defect can only be unfused if $\beta=a\star\alpha$.
3. RDMs in the presence of defects.— The physical situation we consider is that of a circular constant time slice pierced by a topological defect. To be unambiguous, this setup needs to be completed by assigning a proper defect network. In (1) this is achieved by closing the defect on itself through compactification – at the price of problems ((P1)) and ((P2)). The route we shall pursue here instead, is to have the defect emanate from a twist field $\phi$ prepared in the distant past, see figure LABEL:fig:RDM_state. Two significant modifications compared to $\rho_{A}^{old}$ arise:
@@
It is the goal of this letter to demonstrate that (M1) & (M2) are feature, not a bug, and in fact essential in reproducing the numerical observations in the Ising model Roy_2022. We now proceed to construct RDMs based on these modifications.
@@
The EE associated with $(i)$ or $(iii)$ accounts for quantum correlations in an impurity-doped system $A$ or $B$, respectively. They are called “symmetric entropy” for $A$ or $B$. The EE of $(ii)$, on the other hand, measures quantum correlations across the defect, and is called “interface entropy”.
In order to factorize the Hilbert space $\mathcal{H}\to\mathcal{H}_{{\alpha\beta}}^{A}\otimes\mathcal{H}_{\beta\alpha}^{B}$, we excise a disk surrounding the entangling edges and impose boundary conditions $\alpha,\,\beta$ thereon Ohmori_2015. In case ($ii$), $L$ attaches to the boundary $\alpha$, say. As explained in (2), this is possible only if the fusion rules allow it and a-priori requires the choice of another boundary condition $\alpha^{\prime}\in L\star\alpha$, see figure LABEL:fig:RDM_OT. An RDM is constructed out of this state as in figure LABEL:fig:RDM_spherestrip. Finally, a conformal map allows us to depict RDMs as strips,
@@
up to a normalization enforcing ${\rm Tr}\rho^{\phi}_{{\alpha\beta}}=1$. The strip’s height is 1 and its width $\frac{1}{\pi}|\log\ell(R)/\epsilon|$, with $\ell(R)=2\sin(\pi R)$, is controlled by the interval size $R$ defined in figure LABEL:fig:RDM_OT. The punctures are a distance $R$ apart and entanglement time runs from bottom to top. $\rho^{\phi,S}_{\alpha\beta}$ belongs to the “symmetric” case $(i)$ and $\rho^{\phi,I}_{\alpha\beta}$ to the “interface” case $(ii)$. Replica manifolds for case $(iii)$ are equivalent to $\rho_{\alpha\beta}^{\phi,S}$ with $R\to 1-R$ and are not treated separately here. We stress that the RDMs (3) are not distinguished by the conformal weights $h_{\phi}$ and $\bar{h}_{\phi}$ of the twisted primary, but only by the choice of the defect network. This proves to be crucial, especially when dealing with entanglement in twisted states associated with non-invertible defect lines. All details and our conventions for the factorization procedure are found in Northe:2025qcv.
Note that our framework manifestly avoids the problems ((P1)) and ((P2)). Indeed, the topology of the interval $A$ stays intact. Furthermore, the RDMs (3) act on a chiral Hilbert space $\mathcal{H}_{\alpha\beta}^{A}$, which provides the entanglement spectrum.
@@
where $\mathsf{c}$ is the central charge of the CFT, $\mathsf{g}_{\alpha}=\langle 0\lVert{\alpha}\rangle\!\rangle$ is the Affleck-Ludwig boundary entropy of the boundary condition $\alpha$ and $\mathcal{F}_{n}^{\phi}(R)$ results from $\phi^{\times 2n}$ correlators, e.g. for the vacuum field, $\mathcal{F}_{n}^{\mathbbm{1}}(R)=1$. While a rescaling of the cutoff $\epsilon$ can absorb the $\mathsf{g}_{\alpha}\mathsf{g}_{b}$ term, it can always be recovered by comparing entropies of distinct $n$. Dots stand for vanishing contributions in the limit $\epsilon\to 0$.
We now carry on to reproduce (4) from the RDMs (3) of twisted primary states in the Ising CFT, first for grouplike defects, and thereafter for duality defects. In both cases we match our findings to the numeric results of Roy_2022, thereby validating the modifications ((M1)) & ((M2)).
@@
4. Matching EE for grouplike defects.— Numeric modeling of the symmetric and interface entropy results in the following ground state EEs for the grouplike defect $D_{\varepsilon}$ Roy_2022
@@
On the left, the only choice for $\alpha^{\prime}$ is $\alpha\star\varepsilon$, because grouplike defects map elementary boundaries into themselves. Hence, the defect can be trivially unfused from the boundary. In the first equality, this is done partially, and from the entire left boundary in the second equality, leaving $\alpha\star\varepsilon$. Using $D_{\varepsilon}^{2}=\mathbbm{1}$, the defect line is moved onto the field insertions.
At the price of dividing by $\mathsf{d}_{\sigma}$, we nucleate a $D_{\sigma}$ loop in the second and fourth image of (6). Moving it past the field insertions, $\mu$ turns to $\sigma$, shown exemplarily starting from the second image,
@@
Due to $\mathcal{F}_{n}^{\sigma}(R=1)$ Ibáñez_Berganza_2012, the observation (5) is theoretically confirmed and extended to any $n$. All details are found in the supplemental material.
@@
where $f(R)$ is a known function Klich:2015ina. Its specific form is not necessary for us, only its asymptotic values $f(R)\approx\frac{\log 2}{2}$ for $R\to 1$ and $f(R)\approx 0$ for $R\to 0$. Our framework breaks down at $R=0$ or $1$, since the usual conformal transformations to the replica manifold become singular. So we are not able to test the contribution $\frac{\delta_{R,1}}{2}\log 2$ which would however stem from completing the EE to a regular entropy.
@@
The non-invertible nature of $D_{\sigma}$ makes it difficult to find a fruitful network manipulation relating replica manifolds of $\rho_{\alpha\beta}^{\nu,S}$ and $\rho_{\alpha\beta}^{\nu,I}$ similar to (7). The strategy we pursue is thus indirect: we show that, for a specific choice of intermediary boundary condition $\alpha^{\prime}$ in (3), $S_{n}(\rho_{\alpha\beta}^{\nu,S})$ and $S_{n}(\rho_{\alpha\beta}^{\nu,I})$ are related by $R\to 1-R$, and furthermore that they reproduce the asymptotic values of $f(R)$ for small (large) intervals $R\ll 1$ $(1-R\ll 1)$.
@@
For the interface entropy, we face the important subtlety that one can potentially chose different intermediate boundary conditions $\alpha^{\prime}\in\sigma\star\alpha$ in the RDM $\rho^{\nu,I}_{{\alpha\beta}}$. We claim that the most natural choice, which reproduces the numerical results, is the proper fusion $\alpha^{\prime}=\sigma\star\alpha$. Notice that if $\alpha=\sigma$, then $\alpha^{\prime}=\mathbbm{1}+\varepsilon$ is a non-elementary boundary condition, which poses no problem, however. With this choice, the two $D_{\sigma}$ lines can be unfused from the boundary similarly to the first move in (6). The resulting RDM looks like the right-hand side of (10) but with 1R1-R replaced by $R$. This shows $S_{n}(\rho_{\alpha\beta}^{\nu,S},R)=S_{n}(\rho_{\alpha\beta}^{\nu,I},1-R)$, as claimed. We remark that this type of argument does not apply in the conventional twist field picture of entanglement Calabrese_2009, since detaching the defect from the entangling edge must be justified by alternate means.
To complete our argument, we now match the asymptotic values of the EE in the symmetric case. As $R\to 0$ the two $\nu$ punctures approach each other head first. We can use the OPE $\nu(z)\nu(w)=\frac{\mathbbm{1}_{\sigma}}{(z-w)^{1/8}}+\cdots$, where |zw|=R|z-w|=R and the dots vanish as $z\to w$. Because the defect tails connect into a continuous line, $\mathbbm{1}_{\sigma}$ is the identity field on $D_{\sigma}$. To leading order in $1/R$, (10) simplifies to
@@
where in the last step we fused the line to the right boundary and included normalizations $\cal N,N^{\prime}$. The overall prefactor $R^{-1/8}$ cancels out due to the normalization and does not contribute to the Rényi entropy. This shows that $S_{n}(\rho^{\sigma,S}_{{\alpha\beta}})\stackrel{{\scriptstyle R\ll 1}}{{\approx}}S_{n}(\rho^{\mathbbm{1}}_{\alpha,\sigma\star\beta})$. Using (4) and the relation $\mathsf{g}_{a\star b}=d_{a}\mathsf{g}_{b}$, together with $d_{\sigma}=\sqrt{2}$ yields the expected result $S_{n}(\rho^{\sigma,S}_{{\alpha\beta}})\stackrel{{\scriptstyle R\ll 1}}{{\approx}}S_{n}(\rho^{\mathbbm{1}}_{{\alpha\beta}})+\frac{\log 2}{2}$ for any choice of boundary conditions $\alpha,\beta$.
In the opposite regime, $R\to 1$, the $\nu$ punctures in the right-hand side of (10) collide tails first. Using the $\nu\nu$ OPE once more, only this time with bulk identity field $\mathbbm{1}$, we immediately obtain the asymptotic behavior $S_{n}(\rho^{\sigma,S}_{{\alpha\beta}})\stackrel{{\scriptstyle 1-R\ll 1}}{{\approx}}S_{n}(\rho^{\mathbbm{1}}_{{\alpha\beta}})$ for any choice of boundary conditions. This concludes our proof, which extends (9) to any replica index $n$. In fact, our arguments are entirely independent of $n$, so that they persist in any analytic continuation in $n$.
6. Discussion and Outlook.— In summary, we argued that in order to study entanglement in the presence of defects, the allowed space of states must be enlarged to include twisted sectors, ((M1)). Rényi entropies of twisted excitations ((M2)) then reproduce numerically observed EEs, which we confirmed in the case of the Ising model for its two non-trivial topological defects. This construction manifestly avoids the problems ((P1)) & ((P2)) of (1). Moreover, our arguments are fairly general and apply to rational CFTs with non-diagonal spectrum, given appropriate adaptations in e.g. fusion products.
@@
Our results show that the choice of intermediate boundary condition $\alpha^{\prime}$ in (3) is crucial. Indeed, our analysis rests on the circumstance that the defects can be unfused from the entangling edge and moved into the region $B$. This is possible for the “natural” choice $\alpha^{\prime}=\alpha\star a$, which can be non-elementary. In the grouplike case, $a=\varepsilon$, this is the only choice. In the duality case $a=\sigma$, on the other hand, any $\alpha^{\prime}\in\alpha\star\sigma$ may be chosen, and the defect cannot be unfused from the boundary a priori. One can furthermore investigate entanglement in junction fields, i.e. fields sitting at junctions of an arbitrary number of defects. This is left to future work.

### v2

@@ Abstract
Present theoretical predictions for the entanglement entropy through topological defects are violated by numerical simulations. In order to resolve this, we introduce a paradigm shift in the preparation of reduced density matrices in the presence of topological defects, and emphasize the role of defect networks with which they can be dressed. We consider the cases of grouplike and duality defects in detail for the Ising model, and find agreement with all numerically found entanglement entropies. Since our construction functions at the level of reduced density matrices, it accounts for topological defects beyond the entanglement entropy to other entanglement measures.
1. Introduction.— Defects and interfaces appear ubiquitously in nature, from impurities in materials AffleckGFactor over quantum dots sierra2014strongly and tensor networks Hauru:2015abi all the way to D-branes in string theory Polchinski:1998rr. The last decade has, moreover, witnessed a reformulation of our understanding of symmetries to what is now known as generalized symmetry Shao:2023gho – a development driven mainly by topological defects. These are completely energy-transmissive and can therefore be deformed at will, until fields or boundaries are encountered, where topological defects act in various interesting ways. In addition to implementing conventional group actions, topological defects can implement, for instance, order-disorder dualities Frohlich_2007, most famously Kramers-Wannier duality in the Ising model Frohlich:2004ef.
@@
The effect of topological defects, in contrast, on entanglement is more nuanced. They appear for instance naturally as a tool for symmetry resolution of entanglement spectra Belin_2013; Goldstein:2017bua; Zhao_2021; Weisenberger_2021; Zhao:2022wnp; northe2023entanglement; Kusuki:2023bsp; Choi:2024wfm; Kusuki:2024gss; Heymann_2025; Northe:2025qcv; benini2025entanglementasymmetryhighernoninvertible; Bhattacharyya:2025tmg; Saura-Bastida:2024yye; Das:2024qdx; magan2021proof; capizzi2022symmetry; Capizzi:2022igy, where the defects are imposed manually. Situations where the defect forms part of the physical system, crossing in particular through one entangling edge, were studied early on Sakai_2008; Brehm:2015lja; Brehm_2016; Brehm:2017kuz; Gutperle_2016 in high-energy theory. Their proposed framework and results are widely accepted beyond high energy physics in quantum information and condensed matter Eisler:2012xry; Wen:2017smb; Afxonidis:2024gne; Gutperle:2017enx; gutperle2024noteentanglemententropytopological; capizzi2023domain; Capizzi:2023mly; Capizzi:2023vsz. Said framework Sakai_2008; Brehm_2016; Gutperle_2016 predicts a subleading contribution to the EE, which, however, has recently been seen to be contradicted in ab initio numerical simulations of the Ising model Roy_2022; Rogerson:2022yim, the 3-state Potts model Sinha:2023hum and the Luttinger liquid Roy:2023wer.
Subleading contributions encode the intricate interplay between boundaries, defects, but also field excitations Ibáñez_Berganza_2012 and the entanglement spectrum Li:2008kda; Qi:2012ngv. The latter is the eigenspectrum of the reduced density matrix (RDM) and captures therefore much more information than the EE. It is thus important to have a formalism capable of predicting the effect of topological defects on quantum correlations correctly. The purpose of this letter is to fill this gap.
It is useful to begin with a sketch of the prevailing theoretical framework Sakai_2008. Consider an interface $L$ separating the right-half from the left-half plane, and denominate the positive real line as entangling interval $A$. A reduced density matrix $\rho_{A}^{old}$ is prepared by identifying two concentric circles centered at the entangling edge $\partial A$, one with radius $\epsilon\ll 1$, and one with radius $1/\epsilon$,
@@
In this way the interval $A$ is compactified onto a circle, and the RDM is a defect-dressed cylinder in which entanglement time runs from bottom to top. Because the identification splits the defect in two, it acts twice on the subregion Hilbert space $\mathcal{H}_{A}$. Stacking $n$ such RDMs and tracing over $\mathcal{H}_{A}$ leads to a torus. Although, this framework produces the correct leading order in the EE, two problems are apparent:
@@
The RDM (1) predicts subleading terms in the EE Brehm_2016; Gutperle_2016, which depend explicitly on two defects $L$ instead of one, due to ((P1)), apply to a compact subregion, instead of a line segment, again due to ((P1)), and draw on its non-chiral entanglement spectrum, due to ((P2)). Precisely these predictions are violated in the simulations Roy_2022; Rogerson:2022yim; Sinha:2023hum; Roy:2023wer. A further element of contention is that the simulated subleading terms depend generically on the subregion size, which is not mirrored by the theory of Sakai_2008.
In this letter, we construct RDMs for an interval $A$, which do not suffer from these problems and reproduce the simulations in the Ising model Roy_2022. Besides having $L$ pierce the entangling edge $\partial A$, we also discuss the situations where $L$ crosses inside and outside of $A$. Our guiding principle is to consider proper choices of defect network that complete the defect piercing the constant time slice. Ultimately, this leads us to introduce twist fields and study their quantum correlations. We work directly at the level of RDMs, so that our formalism’s utility extends beyond the EE to the investigation of other entanglement measures.
After briefly recapitulating boundaries and defects, we construct our RDMs and match their Rényi entropies to numerically obtained EEs in the Ising model Roy_2022.
@@
Note that in the second picture the defect can only be unfused if $\beta=a\star\alpha$. Lastly, the Affleck-Ludwig boundary entropy $\mathsf{g}_{\alpha}=\langle 0\lVert{\alpha}\rangle\!\rangle$ AffleckGFactor changes upon application of a Verlinde line to $\mathsf{g}_{a\star\alpha}=\langle{0}|D_{a}\lVert{\alpha}\rangle\!\rangle=\mathsf{d}_{a}\mathsf{g}_{\alpha}$, where $\mathsf{d}_{a}$ is the quantum dimension.
3. RDMs in the presence of defects.— The physical situation we consider is that of a circular constant time slice pierced by a topological defect. To be unambiguous, this setup needs to be completed by assigning a proper defect network. In (1) this is achieved by closing the defect on itself through compactification – at the price of problems ((P1)) & ((P2)). The route we shall pursue here instead, is to have the defect emanate from a twist field $\phi$ prepared in the distant past, see figure LABEL:fig:RDM_state. Two significant modifications compared to $\rho_{A}^{old}$ arise:
@@
It is the goal of this letter to demonstrate that ((M1)) & ((M2)) are feature, not a bug, and in fact essential in reproducing the numerical observations in the Ising model Roy_2022. We now proceed to construct RDMs based on these modifications.
@@
The EE associated with $(i)$ or $(iii)$ accounts for quantum correlations in an impurity-doped system $A$ or $B$, respectively. They are called “symmetric entropy” for $A$ or $B$. The EE of $(ii)$, on the other hand, measures quantum correlations across the defect, and is called “interface entropy.”
In order to factorize the Hilbert space $\mathcal{H}\to\mathcal{H}_{{\alpha\beta}}^{A}\otimes\mathcal{H}_{\beta\alpha}^{B}$, we excise a disk surrounding the entangling edges and impose boundary conditions $\alpha,\,\beta$ thereon Ohmori_2015. In case ($ii$), $L$ attaches to the boundary $\alpha$, say. As explained in (2), this is possible only if the fusion rules allow it and a-priori requires the choice of another boundary condition $\alpha^{\prime}\in L\star\alpha$, see figure LABEL:fig:RDM_OT. An RDM is constructed out of this state and its adjoint as in figure LABEL:fig:RDM_spherestrip. Finally, a conformal map allows us to depict RDMs as strips,
@@
up to a normalization enforcing ${\rm Tr}\rho^{\phi}_{{\alpha\beta}}=1$. The strip’s height is 1 and its width $\frac{1}{\pi}|\log\ell(R)/\epsilon|$, with $\ell(R)=2\sin(\pi R)$, is controlled by the interval size $R$ defined in figure LABEL:fig:RDM_OT. The punctures are a distance $R$ apart and entanglement time runs from bottom to top. $\rho^{\phi,S}_{\alpha\beta}$ belongs to the “symmetric” case $(i)$ and $\rho^{\phi,I}_{\alpha\beta}$ to the “interface” case $(ii)$. Replica manifolds for case $(iii)$ are equivalent to $\rho_{\alpha\beta}^{\phi,S}$ with $R\to 1-R$ and are not treated separately here. We stress that the two RDMs in (3) are not distinguished by the conformal weights $h_{\phi}$ and $\bar{h}_{\phi}$ of the twisted primary, but only by the defect network. This proves to be crucial, especially when dealing with entanglement in twisted states associated with non-invertible defect lines. All details and our conventions for the factorization procedure are found in Northe:2025qcv.
Note that our framework manifestly avoids the problems ((P1)) & ((P2)). Indeed, the topology of the interval $A$ and the defect network in the original configuration stay intact. Furthermore, the RDMs (3) act on a chiral Hilbert space $\mathcal{H}_{\alpha\beta}^{A}$, which provides the entanglement spectrum. The main message of our letter is thus that entanglement through defects is captured by the RDMs (3) instead of the RDM (1).
@@
where $\mathsf{c}$ is the central charge of the CFT and $\mathcal{F}_{n}^{\phi}(R)$ results from $\phi^{\times 2n}$ correlators, e.g. for the vacuum field, $\mathcal{F}_{n}^{\mathbbm{1}}(R)=1$. While a rescaling of the cutoff $\epsilon$ can absorb the $\mathsf{g}_{\alpha}\mathsf{g}_{b}$ term, it can always be recovered by comparing entropies of distinct $n$. Dots stand for vanishing contributions in the limit $\epsilon\to 0$.
We now carry on to reproduce (4) from the RDMs (3) of twisted primary states in the Ising CFT, first for grouplike defects, and thereafter for duality defects. In both cases we find agreement with the numeric results of Roy_2022, thereby validating the modifications ((M1)) & ((M2)), and thus the RDMs (3).
@@
4. Matching EE for grouplike defects.— Numeric modeling of the symmetric and interface entropy results in the following ground state symmetric (S) and interface (I) EEs for the grouplike defect $D_{\varepsilon}$ Roy_2022
@@
On the left, the only choice for $\alpha^{\prime}$ is $\alpha\star\varepsilon$, because grouplike defects map elementary boundaries into themselves. Hence, the defect $D_{\varepsilon}$ can be trivially unfused from the boundary. In the first equality, this is done partially, and from the entire left boundary in the second equality, leaving $\alpha\star\varepsilon$. Using $D_{\varepsilon}^{2}=\mathbbm{1}$, the defect line is moved onto the field insertions.
At the price of dividing by the quantum dimension $\mathsf{d}_{\sigma}$, we nucleate a $D_{\sigma}$ loop in the second and fourth image of (6). Moving it past the field insertions, $\mu$ turns to $\sigma$, shown exemplarily starting from the second image,
@@
Due to $\mathcal{F}_{n}^{\sigma}(R)=1$ Ibáñez_Berganza_2012, the observation (5) is theoretically confirmed and extended to any $n$. All details are found in the supplemental material.
@@
where $f(R)$ is a known function Klich:2015ina. Its specific form is not necessary for us, only its asymptotic values $f(R)\approx\frac{\log 2}{2}$ for $R\to 1$ and $f(R)\approx 0$ for $R\to 0$. Our framework breaks down at $R=0$ or $1$, since the usual conformal transformations to the replica manifold become singular. So we are not able to test the contribution $\frac{\delta_{R,1}}{2}\log 2$ which would however stem from completing the EE to a regular entropy. It is important to recall that $S_{1}(\mathbbm{1})$ is invariant under $R\to 1-R$, so that $S_{1}^{S}(D_{\sigma})$ and $S_{1}^{I}(D_{\sigma})$ are exchanged under this swap (away from $R=1$).
@@
The non-invertible nature of $D_{\sigma}$ makes it difficult to find a network manipulation relating replica manifolds of $\rho_{\alpha\beta}^{\nu,S}$ and $\rho_{\alpha\beta}^{\nu,I}$ similar to (7). Furthermore, no bosonized representation for the field $\nu$ is known to us with which we could parallel the analysis in Ibáñez_Berganza_2012. Nevertheless, we provide strong evidence in favor of the RDMs (3): we show that, for a specific choice of intermediary boundary condition $\alpha^{\prime}$ in (3), $S_{n}(\rho_{\alpha\beta}^{\nu,S})$ and $S_{n}(\rho_{\alpha\beta}^{\nu,I})$ are related by $R\to 1-R$, and furthermore that they reproduce the asymptotic values of $f(R)$ for small, $R\ll 1$, and large intervals, $1-R\ll 1$. These three non-trivial matches establish high confidence that the RDMs (3) indeed describe the $D_{\sigma}$-dressed subsystem.
@@
For the interface entropy, we face the important subtlety that one can potentially choose different intermediate boundary conditions $\alpha^{\prime}\in\sigma\star\alpha$ in the RDM $\rho^{\nu,I}_{{\alpha\beta}}$. We claim that the most natural choice, which reproduces the numerical results, is the proper fusion $\alpha^{\prime}=\sigma\star\alpha$. Notice that if $\alpha=\sigma$, then $\alpha^{\prime}=\mathbbm{1}+\varepsilon$ is a non-elementary boundary condition, which poses no problem, however. With this choice, the two $D_{\sigma}$ lines can be unfused from the boundary similarly to the first move in (6). The resulting RDM looks like the right-hand side of (10) but with 1R1-R replaced by $R$. This shows $S_{n}(\rho_{\alpha\beta}^{\nu,S},R)=S_{n}(\rho_{\alpha\beta}^{\nu,I},1-R)$, as claimed. We remark that this type of argument does not apply in the conventional twist field picture of entanglement Calabrese_2009, since detaching the defect from the entangling edge must be justified by alternate means.
To complete our argument, we now match the asymptotic values of the EE in the symmetric case. As $R\to 0$ the two $\nu$ punctures approach each other head first. We can use the OPE $\nu(z)\nu(w)=\frac{\mathbbm{1}_{\sigma}}{(z-w)^{1/8}}+\cdots$, where |zw|=R|z-w|=R and the dots vanish as $z\to w$. Because the defect tails connect into a continuous line, $\mathbbm{1}_{\sigma}$ is the identity field on $D_{\sigma}$. To leading order in $1/R$ (10) simplifies to
@@
where in the last step we fused the line to the right boundary and included normalizations $\cal N,N^{\prime}$. The overall prefactor $R^{-1/8}$ cancels out due to the normalization and does not contribute to the Rényi entropy. This shows that $S_{n}(\rho^{\sigma,S}_{{\alpha\beta}})\stackrel{{\scriptstyle R\ll 1}}{{\approx}}S_{n}(\rho^{\mathbbm{1}}_{\alpha,\sigma\star\beta})$. Using (4) and the relation $\mathsf{g}_{a\star\beta}=d_{a}\mathsf{g}_{\beta}$, together with $d_{\sigma}=\sqrt{2}$ yields the expected result $S_{n}(\rho^{\sigma,S}_{{\alpha\beta}})\stackrel{{\scriptstyle R\ll 1}}{{\approx}}S_{n}(\rho^{\mathbbm{1}}_{{\alpha\beta}})+\frac{\log 2}{2}$ for any choice of boundary conditions $\alpha,\beta$.
In the opposite regime, $R\to 1$, the $\nu$ punctures in the right-hand side of (10) collide tails first. Using the $\nu\nu$ OPE once more, only this time with bulk identity field $\mathbbm{1}$, we immediately obtain the asymptotic behavior $S_{n}(\rho^{\sigma,S}_{{\alpha\beta}})\stackrel{{\scriptstyle 1-R\ll 1}}{{\approx}}S_{n}(\rho^{\mathbbm{1}}_{{\alpha\beta}})$ for any choice of boundary conditions. This concludes our argument, which extends (9) to any replica index $n$. In fact, our reasoning is entirely independent of $n$, so that it persist in any analytic continuation in $n$.
6. Discussion and Outlook.— In summary, we argued that in order to study entanglement in the presence of defects, the allowed space of states must be enlarged to include twisted sectors, ((M1)), and one must study quantum correlations of twisted excitations ((M2)), paying particular attention to the defect network. These modifications lead us to construct our main result, the RDMs (3), with which we reproduce the numerically observed EEs in the Ising model for its non-trivial topological defects. Our construction manifestly avoids the problems ((P1)) & ((P2)) of (1). Moreover, our arguments are fairly general and apply to rational CFTs with non-diagonal spectrum, given appropriate adaptations in e.g. fusion products.
@@
Our results show that the choice of intermediate boundary condition $\alpha^{\prime}$ in (3) is crucial. Indeed, our analysis rests on the circumstance that the defects can be unfused from the entangling edge and moved into the region $B$. This is possible for the “natural” choice $\alpha^{\prime}=\alpha\star a$, which can be non-elementary. In the grouplike case, $a=\varepsilon$, this is the only choice. In the duality case $a=\sigma$, on the other hand, any $\alpha^{\prime}\in\alpha\star\sigma$ may be chosen, and the defect cannot be unfused from the boundary a priori. One can furthermore investigate entanglement in junction fields, i.e. fields sitting at junctions of an arbitrary number of defects. Both options are left to future work.